# Nonparametric Proportional Hazards Model with Differential Regularization Applied to Spatial Survival Data

### v4

@@ 5 Simulation Results
To evaluate the performance of the proposed model, we conducted simulations on a horseshoe-shaped domain $\Omega$ as described in ramsay2002spline and wood2008soap. The true spatial effect is displayed in Figure 2 $(a)$. Note that $\int_{\Omega}h_{0}(\bm{p})d\bm{p}\approx 0.$ We observed independent replicates of $(Y,\delta,\bm{X},\bm{P})=\left(\min(T,C),I(T\leq C),\bm{X},\bm{P}\right),$ where $\bm{X}=(X_{1},X_{2})$ has $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\sim\text{Bernoulli}(0.5)$, and $\bm{P}$ is drawn uniformly from a grid over the domain with spacing 0.025. Conditional on $(\bm{X},\bm{P})$, the event time $T$ follows an exponential distribution with rate parameter $\exp(-4\log(10)+\bm{X}^{\top}\bm{\beta}_{0}+h_{0}(\bm{P})),$ where $\bm{\beta}_{0}=(0.25,-1)^{\top}$, matching the model in (1). The censoring time $C$ is exponentially distributed with rate varying in $\{0.27,2.00\}$, corresponding to approximately $15\%$ and $30\%$ censoring, respectively. The analysis is conducted using sample sizes of $n=500,1000,$ and $2000$.
We construct the mesh using the package of sangalli2021spatial and refine the domain so that the maximum area $\eta$ of the mesh triangles is consistent with the asymptotic analysis, setting $\eta=0.1n^{-0.45}$. We employ first-order FEM basis functions to reduce computational complexity.
@@
$(a)$ True spatial effect $h_{0}$ over the horseshoe-shaped domain $\Omega$.
$(b)$ Sampling design and finite-element mesh for $n=500$ (30% censoring).
$(c)$ Estimated spatial effect $\hat{h}$, averaged over $N=250$ replications.
$(d)$ Average pointwise error surface, $h_{0}-\hat{h}$.
Figure 2: Spatial simulation results: $(a)$ true effect, $(b)$ sampling design, $(c)$ estimated effect, and $(d)$ pointwise error for $n=500$ with 30% censoring.
@@
Table 1: Estimation performance of the proposed spatial PH model compared with the standard PH model and the GAM PH model across different sample size ($n$) and censoring levels. Reported metrics include coefficient bias, mean squared error (MSE) of the regression parameters, $L^{2}$ norm of bias for the spatial effect estimates (only applicable to models with spatial components), and empirical 95% coverage probabilities (CP95) for the regression coefficients in the proposed model.
@@ 6 Empirical Application
In this section, we present two empirical studies that illustrate the practical use of the proposed methodology. The first examines fire brigade response times across Greater London, while the second focuses on crowdsourced seismic data collected during an earthquake in the Campi Flegrei area, Italy. Together, these applications highlight the flexibility of the model across very different domains: urban public safety and real-time earthquake monitoring.
6.1 London Fire Brigade response-time
We illustrate the proposed methodology using the London Fire Brigade response-time dataset, available through the spatsurv package in R. This dataset records the time taken for the first fire engine to arrive at the scene of a dwelling fire across London in 2009 (taylor2017). As in that study, the dataset is augmented with covariates reflecting both spatial and temporal drivers of hazard variability. Harmonic regression terms are included to capture time-of-day periodicity, using the first four pairs of sine–cosine terms: $\{\sin(2\pi i,t/24),\cos(2\pi i,t/24)\},\quad i=1,\dots,4,$ where $t$ denotes the time of day in hours. These covariates provide a flexible approximation of diurnal cycles in fire service response dynamics.
Spatial heterogeneity is introduced via fire station density. Fire station locations were obtained from the fs dataset included in the spatsurv package. Kernel-smoothed intensity of stations was computed and rescaled, providing the covariate fsintens, which captures spatial variation in coverage intensity. The full covariate matrix thus includes $\texttt{fsintens},\{s_{i},c_{i}\}_{i=1}^{4}.$ The final dataset consists of 6,703 uncensored observations.
Following the procedure described in Section 5, we construct a two-dimensional finite element mesh conforming to the administrative boundary of Greater London. The mesh is obtained by refining the boundary geometry and consists of $498$ vertices connected through triangular elements. The resulting mesh is displayed in Figure 3, overlaid with observed fire incident locations.
Figure 3: Fire incident locations coloured by fire engine arrival time (in minutes) across Greater London. Grey lines indicate the triangular finite element mesh, with the thick black outline marking the administrative boundary. The inset shows the location of Greater London within the United Kingdom.
The estimation procedure and smoothing parameter selection followed the same approach as in Section 5. Table 2 reports the estimated regression coefficients and associated standard errors. The results indicate that fire station intensity (fsintens) has a strong and highly significant positive association with the hazard of response, confirming that denser station coverage increases the likelihood of faster arrivals.
Turning to the harmonic regressors, we see clear evidence of diurnal structure in the first two sine terms. Both $s_{1}$ and $s_{2}$ have negative coefficients and are statistically significant at the 5% level, indicating systematic within-day variation in the hazard. The corresponding cosine terms show different behaviour: $c_{2}$ is positive and significant at the same level, suggesting an additional phase shift in the daily cycle, whereas $c_{1}$ is small and not significant. Higher-order harmonics ($s_{3},c_{3}$ and $s_{4},c_{4}$) are generally weaker in magnitude and mostly not significant at the 5% level, indicating that most of the time-of-day effect is captured by the lower-frequency components.
@@ 6.1 London Fire Brigade response-time
fsintens $0.077$ $0.008$ <0.001<0.001
$s_{1}$ 0.050-0.050 $0.021$ $0.016$
$c_{1}$ $0.002$ $0.019$ $0.931$
$s_{2}$ 0.166-0.166 $0.020$ <0.001<0.001
$c_{2}$ $0.043$ $0.020$ $0.028$
$s_{3}$ 0.002-0.002 $0.020$ $0.919$
$c_{3}$ $0.031$ $0.019$ $0.109$
$s_{4}$ 0.031-0.031 $0.018$ $0.089$
$c_{4}$ 0.034-0.034 $0.018$ $0.067$
Table 2: Estimated regression coefficients, standard errors, and p-values for the London Fire Brigade response-time dataset using the proposed spatial PH model.
The estimated nonparametric spatial effect in Figure 4 $(a)$ reveals residual heterogeneity across Greater London after adjusting for covariates. Elevated values are concentrated in the eastern boroughs, particularly toward the city perimeter, while lower values appear in the south-west and west. These patterns suggest that unobserved spatially structured factors—such as traffic congestion or deployment practices—may influence response times beyond station density and diurnal variation.
Figure 4 $(b)$ shows the estimated relative hazard at 12:00 PM. The surface reveals pronounced spatial heterogeneity, with higher relative hazard in central and eastern areas—indicating faster expected arrival times—and lower values in peripheral southern and western boroughs. These patterns mirror the spatial field and confirm the presence of meaningful residual spatial variation after accounting for the covariates.
$(a)$ Estimated spatial effect $\hat{h}(\bm{p})$. Higher values indicate greater hazard (faster expected arrivals), lower values the opposite.
$(b)$ Estimated relative risk $\exp(\bm{x}^{\top}\hat{\bm{\beta}}+\hat{h}(\bm{p}))$ at 12:00 PM. Warmer colors denote faster expected arrivals.
Figure 4: Estimated spatial effect $(a)$ and relative risk $(b)$ for fire engine response times across Greater London.
Overall, the empirical analysis highlights the practical value of the proposed spatial proportional hazards model for understanding emergency response dynamics. The regression results underscore the importance of both fire station coverage and diurnal cycles, while the spatial effect uncovers systematic geographic variation not captured by observed covariates. The model achieves a realistic representation of spatial risk across London, offering clear and interpretable insights into the spatial distribution of emergency response performance.
@@ 6.2 Crowdsourced seismic data analysis
Earthquake Network (EQN) is a citizen science initiative for real time earthquake monitoring using crowdsourced smartphones (finazzi2016). When a smartphone detects shaking, it sends a signal (called trigger) to a central server. Based on all the triggers received, the server decides whether an earthquake is occurring. If an earthquake is detected, an alert containing a warning about the incoming seismic wave is sent to the population. Earthquake detection also represents a censoring event, meaning any subsequent triggers are ignored.
Here, we examine the EQN triggering times collected during the magnitude 3.1 earthquake occurred on 18 February 2025 at 02:22:19 UTC. The epicentre was located in the Campi Flegrei area (Italy) at a depth of 2 km (see Figure 5);
@@
As before, we construct a two-dimensional finite element mesh that conforms to the irregular boundary of cost of Campi Flegrei and the distance limit from the epicentre. The mesh is obtained by refining the boundary geometry and consists of $229$ vertices connected through triangular elements. The resulting mesh, along with smartphone locations, is shown in Figure 6 $(a)$.
Figure 5: Map of censored and uncensored observations/smartphones for the 18 February 2025 earthquake in Campi Flegrei (Italy). Grey lines depict the triangular mesh, while the thick black outline marks the administrative boundary. Blue and red dots represent uncensored and censored observations, respectively. The black star denotes the earthquake epicentre. The inset map shows the location of the study area, highlighted in red, within Italy.
The estimation procedure and the the smoothing parameter followed the same approach as in Section 5. The estimated hypocentral distance effect is equal to 1.124-1.124, with standard error $0.102$ and p-value <0.01<0.01. This confirms that, as expected, smartphones located further from the epicentre have a lower hazard of triggering, that is, they tend to trigger later or not at all. On the other hand, Figure 6 $(b)$ displays the estimated spatial effect after accounting for the hypocentral distance. The model uncovers a clear east–west gradient across the Campi Flegrei area, which is consistent with the findings of finazzi2025 who obtained the amplification map for the Campi Flegrei area from the analysis of the smartphone shaking intensity rather than triggering time. Finally, Figure 6 $(c)$ illustrates the estimated spatial variation of the relative hazard for the specific earthquake. The highest hazard values are obviously concentrated around the epicentre. However, due to the spatial effect considered in the model (which describes site amplification), the gradient of the hazard is not radial with respect to the epicentre.
$(a)$ Smartphone triggering times during the 18 February 2025 Campi Flegrei (Italy) earthquake. Concentric circles mark 1 km and 2 km distances from the epicentre.
$(b)$ Estimated spatial effect $\hat{h}(\bm{p})$. Higher values indicate faster smartphone triggering (greater hazard).
$(c)$ Estimated relative risk $\exp(\bm{x}^{\top}\hat{\bm{\beta}}+\hat{h}(\bm{p}))$. Warmer colours denote higher hazard (earlier triggering). The black star marks the epicentre; circles show 1 km and 2 km distances.
Figure 6: Campi Flegrei earthquake results: $(a)$ smartphone triggering times, $(b)$ estimated spatial effect, and $(c)$ relative risk on 18 February 2025.
@@ 7 Conclusions
We introduced a non-parametric proportional hazards model that embeds a smooth spatial effect within the classical partial likelihood framework. By combining a Laplacian-based penalty with finite element sieves, the method handles irregular domains while yielding a tractable concave optimization problem and retaining interpretable regression coefficients.
From a theoretical perspective, we established the identification of the model, the consistency of the sieve estimator and the asymptotic normality of the parametric component, thus enabling standard inference. Empirically, simulations show that our modelling approach allows the latent spatial effect to be reconstructed, reduces parameter error relative to standard PH models and produces high-resolution, interpretable risk maps. This approach has been applied to two case studies, in which identifying the spatial effect has provided insight into the data generation mechanism.
While our focus is spatial survival analysis, the framework extends naturally to areal data and more broadly to settings where non-parametric effects must be modelled within a proportional hazards structure.
@@ Appendix C Proof of Theorem 2
The result follows from Theorems 3.1–3.2 of chen2007large for sieve M–estimation, upon verifying, identification and local curvature (Condition 3.1), sieve approximation properties (Condition 3.2), continuity and compactness (Conditions 3.3–3.4), uniform convergence over the sieves (Condition 3.5) and stochastic regularity (Conditions 3.6–3.8), that we state and verify below. Regarding the rate $\epsilon_{n}$, consider that the stochastic term is
@@
We now state and verify Conditions 3.1–3.8 of chen2007large.
Condition 3.1 $(i)$ $Q(\theta_{0})>$-$\infty$, and if $Q(\theta_{0})=+\infty$ then $Q(\theta)<+\infty$ for all $\theta\in\Theta_{k}\setminus\{\theta_{0}\}$ for all $k\geq 1$; $(ii)$ there are a nonincreasing positive function $\delta(\cdot)$ and a positive function $g(\cdot)$ such that for all $\varepsilon>0$ and for all $k\geq 1$,
@@
Proof of Condition 3.1(i)
The value $Q(\theta_{0})$ is finite because $\mathbb{E}\bigl|\bm{X}^{\!\top}\bm{\beta}_{0}+h_{0}(\bm{P})\bigr|<\infty$ and $\mathbb{E}|\log s^{(0)}(\bm{\beta}_{0},h_{0},Y)|<\infty$ by the same moment bounds used in Lemma 3.
Proof of Condition 3.1(ii)
Let $\theta=(\bm{\beta},h)$ be an arbitrary point in the parameter space $\Theta$, and let $\theta_{0}=(\bm{\beta}_{0},h_{0})$ be the true parameter value. We define the perturbation $\xi=\theta-\theta_{0}=(\bm{\beta}-\bm{\beta}_{0},h-h_{0})=(\bm{u},g)$.
@@
Finally, since this bound was established on the full parameter space $\Theta$, it holds automatically for any subset, including the sieve spaces $\Theta_{k}$. Thus, we can set the nonincreasing function $\delta(k)\equiv 1$, which completes the proof.
Condition 3.2 Let the sieve parameter spaces be
@@
where $\mathcal{B}\subset\mathbb{R}^{b}$ is the fixed compact set from Assumption 1 and $\mathcal{H}_{k}\subset\mathcal{H}$ are the finite–element sub-spaces constructed in Section 3.2. Then Condition 3.2 holds:
  1. (i)
    $\Theta_{n}\subset\Theta_{n+1}\subset\Theta$ for every $n\geq 1$;
  2. (ii)
    for the true parameter $\theta_{0}=(\bm{\beta}_{0},h_{0})\in\Theta$ there exists a sequence $\pi_{n}\theta_{0}\in\Theta_{k}$ such that $d(\theta_{0},\pi_{n}\theta_{0})\to 0$ as $k\to\infty$.
Proof of Condtion 3.2(i)
The coefficient space $\mathcal{B}$ is independent of $n$, hence $\mathcal{B}\subset\mathcal{B}$ trivially. For the functional part, our triangulations satisfy $\mathcal{T}_{\eta(n+1)}$ is a refinement of $\mathcal{T}_{\eta(n)}$ with $\eta(n+1)<\eta(n)$; consequently every piecewise-linear function in $\mathcal{H}_{n}$ is also contained in $\mathcal{H}_{k+1}$. Lemma 1 implies that $\Theta_{n}\subset\Theta_{n+1}$.
@@
Condition 3.3 $(i)$ For every fixed $n\geq 1$ the mapping $Q(\theta)$ is upper semicontinuous on $\Theta_{n}$ under $d$; $(ii)$ $\bigl|Q(\theta_{0})-Q(\pi_{n}\theta_{0})\bigr|=o\!\bigl(\delta\!\bigl(k(n)\bigr)\bigr)$ for any sequence $k(n)\uparrow\infty$.
Proof of condition 3.3(i)
Lemma 3 established that $Q$ is (everywhere) continuous on the full space $\Theta$. Continuity clearly implies both upper and lower semicontinuity, so the restriction of $Q$ to any subset – in particular to each finite-dimensional, closed set $\Theta_{n}$ – is automatically upper semicontinuous.
Proof of condition 3.3(ii)
We have shown in (14) that $d\bigl(\theta_{0},\pi_{k}\theta_{0}\bigr)\!\to 0$ as $n\to\infty$. Because $Q$ is continuous (Lemma 3),
@@
which is the assertion, as we set $\delta(k(n))=1$.
Condition 3.4 For every $n\geq 1$ the sieve space $\Theta_{n}=\mathcal{B}\times\mathcal{H}_{n},$ is compact under the metric $d(\cdot,\cdot)$.
Proof of Condition 3.4
By assumption $\mathcal{B}\subset\mathbb{R}^{b}$ is compact. The fact that $\mathcal{H}_{n}$ is compact under the $\|\cdot\|_{\infty}$ is shown in Lemma 2. The Cartesian product of two compact sets is compact, and the metric $d$ is simply the sum of the metrics on the two factors. Thus $\Theta_{n}$ is compact under $d$.
Condition 3.5 $(i)$ for every fixed $k\geq 1$, $\sup_{\theta\in\Theta_{k}}\bigl|\hat{Q}_{n}(\theta)-Q(\theta)\bigr|\xrightarrow{p}0$; $(ii)$ along any sequence $k=k(n)\to\infty$ satisfying $\sup_{\theta\in\Theta_{k(n)}}\bigl|\hat{Q}_{n}(\theta)-Q(\theta)\bigr|=o_{p}(\delta(k))$.
Proof of Condition 3.5(i)
For $Z=(Y,\delta,\bm{X},\bm{P})$ define

### v5

@@ 5 Simulation Results
To evaluate the performance of the proposed model, we conducted simulations on a horseshoe-shaped domain $\Omega$ as described in Ramsay (2002) and Wood et al. (2008). The true spatial effect is displayed in Figure 2 $(a)$. Note that $\int_{\Omega}h_{0}(\bm{p})d\bm{p}\approx 0.$ We observed independent replicates of $(Y,\delta,\bm{X},\bm{P})=\left(\min(T,C),I(T\leq C),\bm{X},\bm{P}\right),$ where $\bm{X}=(X_{1},X_{2})$ has $X_{1}\sim\mathcal{N}(0,1)$ and $X_{2}\sim\text{Bernoulli}(0.5)$, and $\bm{P}$ is drawn uniformly from a grid over the domain with spacing 0.025. Conditional on $(\bm{X},\bm{P})$, the event time $T$ follows an exponential distribution with rate parameter $\exp(-4\log(10)+\bm{X}^{\top}\bm{\beta}_{0}+h_{0}(\bm{P})),$ where $\bm{\beta}_{0}=(0.25,-1)^{\top}$, matching the model in (1). The censoring time $C$ is exponentially distributed with rate varying in $\{0.27,2.00\}$, corresponding to approximately $15\%$ and $30\%$ censoring, respectively. The analysis is conducted using sample sizes of $n=500,1000,$ and $2000$.
We construct the mesh using the package of Sangalli (2021) and refine the domain so that the maximum area $\eta$ of the mesh triangles is consistent with the asymptotic analysis, setting $\eta=0.1n^{-0.45}$. We employ first-order FEM basis functions to reduce computational complexity.
@@
$(a)$
$(b)$
$(c)$
$(d)$
Figure 2: Spatial simulation results: $(a)$ true spatial effect $h_{0}$ over the horseshoe-shaped domain $\Omega$; $(b)$ sampling design and finite-element mesh for $n=500$ and 30% censoring; $(c)$ estimated spatial effect $\hat{h}$, averaged over $N=250$ replications; $(d)$ average pointwise error surface, $h_{0}-\hat{h}$ for $n=500$ with 30% censoring.
@@
Table 1: Estimation performance of the proposed spatial PH model compared with the standard PH model and the GAM PH model across different sample size ($n$) and censoring levels. Reported metrics include coefficient bias, mean squared error (MSE) of the regression parameters, $L_{2}$ norm of bias for the spatial effect estimates (only applicable to models with spatial components), and empirical 95% coverage probabilities (CP95) for the regression coefficients in the proposed model.
@@ 6 Empirical Application
In this section, we present two empirical studies that illustrate the practical use of the proposed methodology. The first examines emergency medical service (EMS) response times for ambulance units in the city of San Francisco, while the second focuses on crowdsourced seismic data collected during an earthquake in the Campi Flegrei area, Italy. Together, these applications highlight the flexibility of the model across very different domains: urban public safety and real-time earthquake monitoring. Both cases demonstrate the model’s ability to handle complex spatial domains—ranging from the intricate coastline of a peninsula to the volcanic topography of a caldera—while adjusting for high-dimensional temporal and environmental covariates.
6.1 San Francisco EMS Ambulance response-time
The proposed methodology is illustrated using San Francisco Division of Emergency Medical Services (EMS) dispatch data of February 2025, comprising 7,890 different interventions. Each intervention records a sequence of chronological milestones, including the time the call was received, the unit dispatch time, and the start of the unit’s response. The event of interest is the duration until the first unit’s arrival on scene, which is subject to right-censoring originating from two distinct sources. First, operational censoring occurs when a unit is diverted to a higher-priority incident or cancelled before reaching the scene. In such instances, the arrival timestamp is missing, and the censoring time is defined as the duration from the initial call to the last recorded operational activity. Second, administrative censoring was applied to limit the influence of extreme outliers by capping all response times at 60 minutes. For reference, the 99th percentile of response time is approximately 37 minutes. Any intervention exceeding 60 minutes was therefore treated as right-censored at that threshold.
The model is adjusted for several environmental and temporal covariates that might influence medical unit mobility. Weather data were sourced from the Open-Meteo archive, and is relative to the average coordinates of the observed locations (37.77 N, 122.42 W). Weather data includes temperature (temp) measured in degrees Celsius, hourly precipitation (precip) in millimeters, and wind speed (wind) in kilometers per hour. Daily periodicity is captured via harmonic terms $s_{k},c_{k}$ for $k=1,2$, representing 24-hour and 12-hour cycles derived from the decimal hour of the call. We also include a binary indicator for high-priority emergency calls (is_emergency), and a holiday indicator (is_holiday) based on the New York Stock Exchange (NYSE) calendar. The latter serves as a proxy for urban traffic dynamics, as NYSE holidays align with the closure of San Francisco’s financial district and school systems, significantly reducing road congestion.
In order to accommodate San Francisco’s complex coastal geometry and internal water bodies, we define the spatial domain of interest as the city’s geographic boundaries where major water features, such as Lake Merced, are excluded, while Treasure Island is retained. The resulting domain is shown in Figure 3 (a), displayed in the context of the state of California. This domain is then discretized using a two-dimensional finite element mesh generated via constrained Delaunay refinement. The resulting triangulation consists of 468 nodes. Figure 3 (b) displays the resulting mesh along with the spatial distribution of EMS incidents, where point colours indicate observed response times in minutes.
$(a)$
$(b)$
Figure 3: Spatial domain and data for San Francisco in February 2025: $(a)$ location of the San Francisco study area within California; $(b)$ finite element mesh and EMS response-time observations, with points indicating incident locations colored according to observed time.
We apply the proposed method, selecting the smoothing parameter $\lambda$ as in the simulation study. The estimated regression coefficients $\hat{\beta}$ are reported in Table 2. The reported numerical values correspond to the estimated coefficients, while the relative $p$-values are computed using standard errors obtained from the inverse Hessian of the penalized log-likelihood.
We observe that the indicator variable is_emergency is the most dominant predictor, with an estimated coefficient of 1.964 $(p<0.001)$. This implies a substantially higher arrival hazard for emergency-status calls. The coefficient for precip is 0.035 and is marginally significant $(p=0.064)$, suggesting a modest increase in the hazard under precipitation. Furthermore, the 12-hour harmonic component $s_{2}$ is statistically significant $(p=0.030)$, indicating a semi-diurnal rhythm in EMS efficiency, possibly associated with peak demand periods or shift changes. The estimated effects of wind, temp, and is_holiday are instead not statistically different from zero at the 10% significance level.
@@ 6.1 San Francisco EMS Ambulance response-time
is_mergency $1.964$ $0.028$ <0.001<0.001
temp 0.002-0.002 $0.005$ $0.685$
precip $0.035$ $0.019$ $0.064$
wind 0.002-0.002 $0.002$ $0.300$
is_holiday $0.038$ $0.026$ $0.150$
$s_{1}$ (24h) 0.007-0.007 $0.018$ $0.680$
$c_{1}$ (24h) 0.022-0.022 $0.021$ $0.305$
$s_{2}$ (12h) 0.037-0.037 $0.017$ $0.030$
$c_{2}$ (12h) $0.004$ $0.017$ $0.814$
Table 2: Estimated covariate effects from the spatial proportional hazards model for San Francisco EMS ambulance response times (February 2025). Estimates are log-hazard ratios; $s_{k}$ and $c_{k}$ denote sine and cosine terms capturing 24-hour ($k=1$) and 12-hour ($k=2$) temporal cycles.
Beyond these global trends, the model uncovers geographic disparities. Figure 4 illustrates the estimated spatial function $\hat{h}(\bm{p})$, where lower values indicate weaker response efficiency. These variations persist after adjusting for covariates, indicating that structural characteristics of the urban environment influence expected arrival times. Higher arrival hazards (faster expected responses) are concentrated in central and north-eastern areas, whereas south-western and south-eastern neighbourhoods exhibit lower hazards. Treasure Island displays a clear north–south gradient, with relatively higher hazards in the north and lower hazards in the south. These disparities likely reflect differences in road-network configuration, station proximity, topographical constraints, and localised access conditions. The model thus provides a representation of EMS risk, offering insights into the geographic and temporal distribution of EMS arrival performance across the urban landscape.
Figure 4: Estimated spatial effect $\hat{h}(\bm{p})$, where higher values correspond to faster expected emergency unit arrivals.
@@ 6.2 Crowdsourced seismic data analysis
Earthquake Network (EQN) is a citizen science initiative for real time earthquake monitoring using crowdsourced smartphones (Finazzi, 2016). When a smartphone detects shaking, it sends a signal (called trigger) to a central server. Based on all the triggers received, the server decides whether an earthquake is occurring. If an earthquake is detected, an alert containing a warning about the incoming seismic wave is sent to the population. Earthquake detection also represents a censoring event, meaning any subsequent triggers are ignored.
Here, we examine the EQN triggering times collected during the magnitude 3.1 earthquake occurred on 18 February 2025 at 02:22:19 UTC. The epicentre was located in the Campi Flegrei area (Italy), as shown in Figure 5(a) at a depth of 2 km.
@@
As before, we construct a two-dimensional finite element mesh that conforms to the irregular boundary of the coastline of Campi Flegrei and the distance limit from the epicentre. The mesh is obtained by refining the boundary geometry and consists of $229$ vertices connected through triangular elements. The resulting mesh, along with smartphone locations colored by censoring status, is shown in Figure 5 (b), while the observed triggering times are illustrated in Figure 5 (c).
$(a)$
$(b)$
$(c)$
Figure 5: Saptial domain and data for Campi Flegrei earthquake data (18 Feb 2025): $(a)$ location of the study area within Italy, $(b)$ triangular mesh with censored (red) and uncensored (blue) observations, and $(c)$ observed smartphone triggering times with distance rings from the epicenter (star) colored according to observed time.
The estimation procedure and the the smoothing parameter followed the same approach as in Section 5. The estimated hypocentral distance effect is equal to 1.124-1.124, with standard error $0.102$ and p-value <0.01<0.01. This confirms that, as expected, smartphones located further from the epicentre have a lower hazard of triggering, that is, they tend to trigger later or not at all. On the other hand, Figure 6 $(a)$ displays the estimated spatial effect after accounting for the hypocentral distance. The model uncovers a clear east–west gradient across the Campi Flegrei area, which is consistent with the findings of Finazzi et al. (2025) who obtained the amplification map for the Campi Flegrei area from the analysis of the smartphone shaking intensity rather than triggering time. Finally, Figure 6 $(b)$ illustrates the estimated spatial variation of the relative hazard for the specific earthquake. The highest hazard values are obviously concentrated around the epicentre. However, due to the spatial effect considered in the model (which describes site amplification), the gradient of the hazard is not radial with respect to the epicentre.
$(a)$
$(b)$
Figure 6: Model results for the Campi Flegrei earthquake: $(a)$ estimated spatial effect $\hat{h}(\bm{p})$, where higher values indicate a greater hazard (faster triggering); $(b)$ estimated relative hazard $\exp(\bm{x}^{\top}\hat{\bm{\beta}}+\hat{h}(\bm{p}))$ at 12:$0$ PM, where warmer colours denote higher values.
@@ 7 Conclusions
We introduced a non-parametric proportional hazards model that incorporates a smooth spatial effect within the classical partial likelihood framework. By combining a Laplacian-based differential penalty with finite element sieve approximations, our approach can handle complex and irregular spatial domains while maintaining a concave optimisation problem and producing interpretable regression coefficients.
From a theoretical perspective, we proved that the model is identifiable, that the sieve estimator is consistent, and that the parametric component is asymptotically normal. These results ensure that standard inferential procedures for the regression coefficients remain valid despite the presence of an infinite-dimensional spatial effect. Empirically, simulation studies demonstrate that the proposed method accurately reconstructs latent spatial variation, reducing bias and mean squared error relative to standard proportional hazards models and yielding high-resolution, interpretable spatial risk surfaces. Two empirical applications further illustrate how explicitly modelling spatial heterogeneity can reveal meaningful structures in the data-generating process that would otherwise remain unobserved.
The proposed methodology, developed for spatial survival analysis, naturally extends to other spatial data structures such as areal and network data, and more broadly to settings requiring non-parametric effects within a proportional hazards framework. The finite element formulation provides a flexible and principled way to represent smooth effects over complex domains, making the approach applicable to a wide range of problems beyond those considered here.
There are several avenues for future research. One important area for further research concerns the goodness-of-fit of the proportional hazards assumption in the presence of a non-parametric spatial effect. Classical diagnostic tools, such as Schoenfeld residuals, are derived under the assumption of purely parametric covariate effects and therefore require careful adaptation in this setting. The development of residual- or score-based diagnostics that explicitly account for the estimated spatial component would enable practitioners to evaluate time-varying effects and identify deviations from proportionality following model fitting. Other avenues to explore include formally quantifying the uncertainty of the spatial effect, selecting the penalty structure based on the data, and extending spatio-temporal survival models to include latent effects that evolve over time.
@@ Appendix C Proof of Theorem 2
The result follows from Theorems 3.1–3.2 of Chen [2007] for sieve M–estimation, upon verifying, identification and local curvature (Condition 3.1), sieve approximation properties (Condition 3.2), continuity and compactness (Conditions 3.3–3.4), uniform convergence over the sieves (Condition 3.5) and stochastic regularity (Conditions 3.6–3.8), that we state and verify below. Regarding the rate $\epsilon_{n}$, consider that the stochastic term is
@@
We now state and verify Conditions 3.1–3.8 of Chen [2007].
Condition 3.1. $(i)$ $Q(\theta_{0})>$-$\infty$, and if $Q(\theta_{0})=+\infty$ then $Q(\theta)<+\infty$ for all $\theta\in\Theta_{k}\setminus\{\theta_{0}\}$ for all $k\geq 1$; $(ii)$ there are a nonincreasing positive function $\delta(\cdot)$ and a positive function $g(\cdot)$ such that for all $\varepsilon>0$ and for all $k\geq 1$,
@@
Proof of Condition 3.1(i): The value $Q(\theta_{0})$ is finite because $\mathbb{E}\bigl|\bm{X}^{\!\top}\bm{\beta}_{0}+h_{0}(\bm{P})\bigr|<\infty$ and $\mathbb{E}|\log s^{(0)}(\bm{\beta}_{0},h_{0},Y)|<\infty$ by the same moment bounds used in Lemma 3. $\square$
Proof of Condition 3.1(ii): Let $\theta=(\bm{\beta},h)$ be an arbitrary point in the parameter space $\Theta$, and let $\theta_{0}=(\bm{\beta}_{0},h_{0})$ be the true parameter value. We define the perturbation $\xi=\theta-\theta_{0}=(\bm{\beta}-\bm{\beta}_{0},h-h_{0})=(\bm{u},g)$.
@@
Finally, since this bound was established on the full parameter space $\Theta$, it holds automatically for any subset, including the sieve spaces $\Theta_{k}$. Thus, we can set the nonincreasing function $\delta(k)\equiv 1$, which completes the proof. $\square$
Let the sieve parameter spaces be
@@
where $\mathcal{B}\subset\mathbb{R}^{b}$ is the fixed compact set from Assumption 1 and $\mathcal{H}_{k}\subset\mathcal{H}$ are the finite–element sub-spaces constructed in Section 3.2. With this notation, we now consider the following condition.
Condition 3.2. (i) $\Theta_{n}\subset\Theta_{n+1}\subset\Theta$ for every $n\geq 1$; (ii) for the true parameter $\theta_{0}=(\bm{\beta}_{0},h_{0})\in\Theta$ there exists a sequence $\pi_{n}\theta_{0}\in\Theta_{k}$ such that $d(\theta_{0},\pi_{n}\theta_{0})\to 0$ as $k\to\infty$.
Proof of Condtion 3.2(i): The coefficient space $\mathcal{B}$ is independent of $n$, hence $\mathcal{B}\subset\mathcal{B}$ trivially. For the functional part, our triangulations satisfy $\mathcal{T}_{\eta(n+1)}$ is a refinement of $\mathcal{T}_{\eta(n)}$ with $\eta(n+1)<\eta(n)$; consequently every piecewise-linear function in $\mathcal{H}_{n}$ is also contained in $\mathcal{H}_{k+1}$. Lemma 1 implies that $\Theta_{n}\subset\Theta_{n+1}$. $\square$
@@
$\square$
Condition 3.3. $(i)$ For every fixed $n\geq 1$ the mapping $Q(\theta)$ is upper semicontinuous on $\Theta_{n}$ under $d$; $(ii)$ $\bigl|Q(\theta_{0})-Q(\pi_{n}\theta_{0})\bigr|=o\!\bigl(\delta\!\bigl(k(n)\bigr)\bigr)$ for any sequence $k(n)\uparrow\infty$.
Proof of condition 3.3(i): Lemma 3 established that $Q$ is (everywhere) continuous on the full space $\Theta$. Continuity clearly implies both upper and lower semicontinuity, so the restriction of $Q$ to any subset – in particular to each finite-dimensional, closed set $\Theta_{n}$ – is automatically upper semicontinuous. $\square$
Proof of condition 3.3(ii): We have shown in (14) that $d\bigl(\theta_{0},\pi_{k}\theta_{0}\bigr)\!\to 0$ as $n\to\infty$. Because $Q$ is continuous (Lemma 3),
@@
which is the assertion, as we set $\delta(k(n))=1$. $\square$
Condition 3.4. For every $n\geq 1$ the sieve space $\Theta_{n}=\mathcal{B}\times\mathcal{H}_{n},$ is compact under the metric $d(\cdot,\cdot)$.
Proof of Condition 3.4: By assumption $\mathcal{B}\subset\mathbb{R}^{b}$ is compact. The fact that $\mathcal{H}_{n}$ is compact under the $\|\cdot\|_{\infty}$ is shown in Lemma 2. The Cartesian product of two compact sets is compact, and the metric $d$ is simply the sum of the metrics on the two factors. Thus $\Theta_{n}$ is compact under $d$. $\square$
Condition 3.5. $(i)$ for every fixed $k\geq 1$, $\sup_{\theta\in\Theta_{k}}\bigl|\hat{Q}_{n}(\theta)-Q(\theta)\bigr|\xrightarrow{p}0$; $(ii)$ along any sequence $k=k(n)\to\infty$ satisfying $\sup_{\theta\in\Theta_{k(n)}}\bigl|\hat{Q}_{n}(\theta)-Q(\theta)\bigr|=o_{p}(\delta(k))$.
Proof of Condition 3.5(i): For $Z=(Y,\delta,\bm{X},\bm{P})$ define